OpenStudy (anonymous):

Find the solutions to the system. y=x^2+8x+2 y=7x+4 (-2,-10) and (1,11) (2,10) and (-1,-11) (-2,-34) and (1,11) or no solutions Please help ive tried everything!! thank you

OpenStudy (freckles):

have you tried setting two two equations equal and solving for x first?

OpenStudy (freckles):

that is you need to first solve: 7x+4=x^2+8x+2

OpenStudy (anonymous):

so far i have 7x=x^2+8x-2

OpenStudy (freckles):

you are one step to having it in the form ax^2+bx+c=0 then you can use quadratic formula or easier to factored the side in which you have ax^2+bx+c for your problem

OpenStudy (freckles):

do you know how to subtract 7x on both sides?

OpenStudy (freckles):

x^2+(8x-7x)-2=0 (is what you have after subtracting 7x on both sides I left the left hand expression for you to simplify

OpenStudy (anonymous):

im sorry im working slow i will simplify it hold on a moment please

OpenStudy (anonymous):

wouldnt it just be x^3-2=0?

OpenStudy (freckles):

well 8-7 isn't 0

OpenStudy (anonymous):

i added the x that was left to the other x^2 and made it x^3

OpenStudy (freckles):

oh I didn't really you but a cube on that x :(

OpenStudy (freckles):

8x-7x is simply 1x not x^3-x^2 or whatever you did

OpenStudy (anonymous):

im sorry i was confused because i thought that putting 1x was wrong because its just the same as x and i thought the x^2 was a like term to x so i added them

OpenStudy (freckles):

1x is the same as x x^2+x doesn't simplify to x^3 (we cannot add x^2 and x because they are not like terms)

OpenStudy (anonymous):

im sorry i was confused, what is the next step?

OpenStudy (freckles):

well you know you have x^2+x-2=0 right?

OpenStudy (freckles):

do you know how to factored x^2+x-2? or do you prefer to use the quadratic formula to solve x^2+x-2=0?

OpenStudy (anonymous):

i dont know either of them very well but i know the quadratic formula probably better

OpenStudy (freckles):

\[ax^2+bx+c=0 \text{ implies you have solutions } x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (freckles):

you need to identify a,b,c as a first step in your equation 1x^2+1x+(-2)=0

OpenStudy (anonymous):

1=a, 1=b,-2=c ?

OpenStudy (freckles):

that is right now you want to plug those numbers into the formula I have above \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \\ x=\frac{-(1)\pm \sqrt{(1)^2-4(1)(-2)}}{2(1)}\]

OpenStudy (freckles):

now simplify that

OpenStudy (freckles):

like I would start with the operations inside the square root

OpenStudy (freckles):

1^2 is the same as 1 can you evaluate 1-4(1)(-2) ?

OpenStudy (anonymous):

is it 6?

OpenStudy (freckles):

1-4(1)(-2) 1-4(-2) 1+8 9 I hope you can see that the inside of the square root doesn't simplify to 6 but it simplifies to 9 --- \[x=\frac{-1 \pm \sqrt{1-4(1)(-2)}}{2(1)} \\ x=\frac{-1 \pm \sqrt{1+8}}{2(1)} \\ x=\frac{-1 \pm \sqrt{9}}{2(1)} \] now that you evaluated the operations inside the square root we can look at the actual square root part what is the square root of the number inside (that is what is the square root of 9?)?

OpenStudy (anonymous):

3?

OpenStudy (freckles):

yes because 3*3 is 9 so you have gotten the square root part out the way so you should have something that looks like this: \[x=\frac{-1 \pm 3}{2}\]

OpenStudy (freckles):

well and I already did the bottom multiplication for you that is I rewrote 2(1) as 2 since 2 times 1 is 2

OpenStudy (freckles):

can you finish simplifying your x values from here?

OpenStudy (anonymous):

im just realy lost to how were eventually going to find the answer to the question

OpenStudy (freckles):

well you have to finish simplifying your x coordinates of your solutions

OpenStudy (anonymous):

oh okay i might need a second hold on please

OpenStudy (anonymous):

(-2,1) ?

OpenStudy (freckles):

do you mean you have x=-2 or x=1 ?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

now find the corresponding y values for each the simplest equation to use is the y=7x+4

OpenStudy (freckles):

for x=-2 you have y=7(-2)+4=? for x=1 you have y=7(1)+4=?

OpenStudy (anonymous):

(-2,-10) (1,11) ?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

thank you so so much!! :)

OpenStudy (freckles):

np

OpenStudy (freckles):

some problems take a little patience especially if you don't know where you are going y=x^2+8x+2 y=7x+4 we wanted to know when these y's were the same x^2+8x+2=7x+4 ( this is why we set up this equality and why we had to solve it ) then after you find the x's you can find the y's by pluggin into either equation

OpenStudy (freckles):

that was just a summary of what we did

OpenStudy (freckles):

anyways goodnight